Find the slopw of a line containing the following points. m(5,3) and (1,6)
Answer by stanbon(75887)

Find the slope of a line containing the following points. m(5,3) and (1,6)
---------------
slope = change in y/change in x
slope = (6-3)/(1-5) = 3/(-4) = -3/4
===================
Cheers,
Stan H.